Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report Page - 20121020

Observers : Patrick El-Hage, Assaf Sternberg. Data reduction team : Nancy and Antonia

All times are in UT

NOTE that due to technical issues, the images were not transferred to wg5off in real time, and we were not able to look at the data. A problem report has been filed, and hopefully the issue will be resolved by tomorrow.

23:45 : EFOSC calibration run using feige110, using 1 slit (seeing measured to be about 1.1).
00:00 : Potential alignment issue with grism. TO looking into it. He has to go to the telescope and fix it manually.
01:00 : EFOSC calibration obs using EG21, using 1 slit for realigned grism setup.
01:15 : EFOSC followup for SN2012dy UBVRi phot and gm11 spec
02:05 : EFOSC followup for LSQ12dyw UBVRi phot and gm13 spec
03:35 : EFOSC followup for SN2009ip UBVRi phot and gm11,gm16 spec
04:20 : EFOSC calibration obs using feige110, 1 slit
04:35 : EFOSC followup LSQ12fhs UBVRi phot and gm11 spec
06:05 : EFOSC followup SN2012ec UBVRi phot and gm11,gm13,gm16 spec
07:30 : EFOSC classification LSQ12fkh
07:45 : EFOSC classification LSQ12fnt
08:10 : EFOSC classification LSQ12fli
08:30 : EFOSC classification LSQ12fmd
Comments